# The Frequency Response of Networks as Open Systems

### v1

@@ Abstract
Many biological, technological, and social systems are effectively networks of interacting individual systems. Typically, these networks are not isolated objects, but interact with their environment through both signals and information that is received by specific nodes with an input function or released to the environment by other nodes with an output function. An important question is whether the structure of different networks, together with the particular selection of input and output nodes, is such that it favors the passing or blocking of such signals. For a given network and a given choice of the input and output nodes, the $\mathcal{H}_{2}$-norm provides a natural and general quantification of the extent to which input signals—whether deterministic or stochastic, periodic or arbitrary—are amplified. We analyze a diverse set of empirical networks and conjecture that many naturally occurring systems—such as food webs, signaling pathways, and gene regulatory circuits—are structurally organized to enhance the passing of signals, facilitating the efficient flow of biomass, information, or regulatory activity. This passing behavior culminates in directed acyclic graphs (DAGs), for which we analytically show that amplification depends on the number and length of input-output pathways, which is consistent with the well-known tendency of naturally emerging networks to approximate DAG structures. In contrast, the structure of engineered systems like power grids appears to be intentionally designed to suppress signal propagation, as the transmitted quantity—voltage phase differences—requires tight control to maintain synchronized operation.
@@ I Introduction
Networked dynamical systems are omnipresent is nature and engineering, ranging from neuronal dynamics to electric power grids. They are made of a multitude of individual interacting units with their own degrees of freedom and intrinsic parameters. The coupling between the degrees of freedom of the individual units typically gives rise to collective behaviors. Usually, these networks are not isolated from their environment, but are subjected to external perturbations, operational changes, and external stimuli. Examples are neuron networks subject to sensory stimuli [1, 2], power grids that both absorb power produced at the generators and deliver power at the loads [3, 4, 5], and networks subject to the effects of environmental noise, such as ecological networks [6]. At the same time, networks produce outputs that affect the environment. Hence, the response of networked systems to external inputs essentially depends on the units chosen as input nodes and output nodes and the structure and characteristics of the network connectivity.
@@
While a large literature has studied control strategies for networks, see e.g., [7, 8, 9, 10], this paper focuses on the response of networked systems to various environmental inputs, not limited to control signals. References [11, 12, 13, 14, 15, 16] have studied the response of a networked system to external stimuli. These studies have primarily focused on graphs with symmetric connectivity or, at most, on asymmetric but normal adjacency matrices. Yet, most real-world networks are highly non-normal, strongly directed, and often exhibit pronounced hierarchical organization—structural features that profoundly shape their dynamical response [17, 18, 19, 20, 21, 22, 23]. Motivated by these properties, our approach does not assume symmetry and instead addresses the response to general input signals, such as sinusoidal, periodic, and stochastic, within a unified framework.
@@
This paper examines how complex networks of coupled dynamical systems respond to environmental interactions, where external signals are received at designated input nodes and measured at specific output nodes [24, 25, 26, 27]. An open network is defined by (i) its internal topology, (ii) the set of input nodes where external stimuli are applied, and (iii) the set of output nodes where the response is measured. As illustrated schematically in Fig. 1, the network structure and its internal connectivity shape how input signals—including sinusoidal, periodic, and noisy—are transformed as they propagate from input to output, with different frequency components affected by amplification, attenuation, or distortion, depending on the structural properties of the network. Such signals are ubiquitous in natural and technological systems—for example, gene regulatory networks responding to circadian rhythms [28], power grids operating at standardized frequencies of 50–60 Hz [29], and neuronal circuits processing signals across a broad frequency range of roughly 0.1–100 Hz [30]—and we emphasize that our framework also applies to noise and arbitrary signals.
@@
Our analysis investigates how the topology of a network and the placement of its input and output nodes shape the amplification or attenuation of external signals. We characterize this behavior through the network’s transfer function, or frequency response, and use the $\mathcal{H}_{2}$-norm as a universal metric that quantifies signal amplification—capturing responses to sinusoidal, periodic, noisy, and general inputs. We establish a direct relationship between the network’s topology, its transfer function, and the controllability Gramian, which serves as a central tool in our study. This framework allows us to assess whether empirical networks tend to amplify (pass) or suppress (block) external inputs for given input-output configurations.
@@
Building on this framework, we perform a broad analysis of empirical networks from different fields, to uncover structural principles governing their signal transformation properties. Key findings are that many biological networks—such as food webs, signaling pathways, and gene regulatory networks—tend to exhibit architectures that enhance the passing of signals, supporting the efficient flow of biomass, information, or regulatory activity. This signal-passing behavior is especially evident in directed acyclic graphs (DAGs), which are known to closely approximate many natural real-world networks—suggesting that such architectures inherently promote signal propagation [18]. Motivated by this, we analytically compute the DC gain for DAG networks, demonstrating that the amplification or attenuation is determined explicitly by the number and length of the input-to-output pathways. In contrast, engineered systems such as power grids—often modeled as undirected (symmetric) graphs—appear to be purposefully designed to suppress signal amplification, reflecting the need for precise regulation of transmitted signals, such as voltage phase differences, to ensure stable and synchronized operation.
@@
FIG. 1: Networks as open systems. The schematic summarizes the goal of the paper in studying how an external signal is altered when it passes through a network. The network is subjected to various input signals from the environment, where only input nodes (in blue) receive them. The response of the network is measured at some output nodes (in orange). A low-frequency sinusoidal input signal may experience a phase shift and amplification, while a high-frequency one may be damped. The network can also suppress noise. The shape of a general output signal may not be the same as its corresponding input signal.
@@ II.1 Frequency Analysis of Networks as Open Systems
Many natural and engineered systems—such as gene networks, power grids, and neural circuits—can be modeled as open networks, where each node evolves through intrinsic dynamics, interacts with neighbors, and communicates with the environment via inputs and outputs. While inherently nonlinear, such systems can often be approximated near equilibrium by a linear time-invariant (LTI) model, enabling powerful analysis through linear control theory. Since many real-world networked systems operate near a stationary equilibrium, we consider a class of nonlinear systems whose linearization around such an equilibrium takes the form $\dot{x}_{i}=-cx_{i}+\sum_{j}A_{ij}x_{j}$, where the scalar $c$ reflects the local stability of the intrinsic dynamics and the adjacency matrix $A$ encodes the network interactions. Throughout this work, we take $c>\max_{i}\Re(\lambda_{i})$, where $\lambda_{i}$ are the eigenvalues of the original adjacency matrix $A$, and apply the spectral shift $A\leftarrow A-c\mathbb{I}$, in order to ensure asymptotic stability by placing all the eigenvalues of $A$ in the left half of the complex plane, while leaving the eigenvectors unaltered.
Note that application of the spectral shift preserves the network’s topology and dynamical response—modifying only the global timescale. This formulation leads to a networked system governed by the standard LTI equations:
@@
where $\boldsymbol{x}(t)\in\mathbb{R}^{N}$ is the network state vector, and $\boldsymbol{u}(t)\in\mathbb{R}^{m}$ is the vector of external input signals acting on the network, reflecting its open nature. The output vector $\boldsymbol{y}(t)\in\mathbb{R}^{p}$ collects signals from selected nodes. The matrices $B$ and $C$ are assumed to have versors as their columns and rows, respectively. That is, if $B_{ij}=1$, then $B_{kj}=0$, $\forall k\neq i$. Similarly for the matrix $C$, if $C_{ij}=1$, then $C_{ik}=0$, $\forall k\neq i$. Our choice of the matrices $B$ and $C$ describes which nodes are input and output nodes, respectively. If $B_{ij}=1$, it means that node $i$ is an input node and receives input signal $j$. Similarly, if $C_{ij}=1$, it means that node $j$ is an output node and transmits output signal $j$. The structure of matrices $B$ and $C$ implies that we can study sets of input nodes and output nodes.
The LTI approximation provides a principled and analytically tractable framework for studying networked systems operating near stable fixed points, during short transients, or under weak perturbations, where linearized dynamics capture the dominant system behavior. This includes many biological and engineered systems functioning under steady-state conditions. Also systems near criticality, such as second-order phase transitions, where linear responses capture essential behavior [31], can be modeled as LTI systems. Even in synchronization problems, governing models in neurodynamics, the LTI framework remains valid when the system is close to the onset of collective oscillations, which typically arise through a Hopf bifurcation occurring at a specific mode (e.g., zero wavenumber or the zero coupling eigenvalue) [32, 33, 34]. In such cases, the dynamics can be linearized around the fixed point prior to bifurcation, yielding time-invariant Jacobians that characterize local behavior. Crucially, we find an important relation with the non-normality of the network adjacency matrix, which is known to induce transient amplification and shape control energy[35] and is a ubiquitous feature of real-world networks. This is especially true for systems with directional flow and hierarchical organization [18, 36], supporting the expectation that many naturally occurring networks exhibit passing behavior when inputs are suitably placed.
In this work, we adopt a frequency-domain approach to characterize the dynamics of networked open systems. Any periodic input can be decomposed into a superposition of sinusoids via Fourier analysis, enabling the study of the system’s frequency response to generic inputs. For the LTI model in Eq. (1), the transfer function is defined as
@@
By evaluating at $s=J\omega$, where $\omega$ is the input frequency and $J=\sqrt{-1}$, we probe the system’s steady-state response to sinusoidal inputs. The output oscillates at the same frequency $\omega$, but with altered magnitude and phase, given by
$\text{Magnitude}=20\log_{10}|G(J\omega)|,\qquad\text{Phase}=\angle G(J\omega).$ (3)
Specifically, if the input signal $u(t)=\sin(\omega t)$, then the output signal $y(t)=\text{Magnitude}\cdot\sin(\omega t+\text{Phase})$. This frequency-based formulation underpins our analysis and enables principled insights into how the network structure shapes signal transmission and network performance.
An important case of the frequency response analysis arises in the low-frequency limit, which corresponds to the case of slowly varying or constant inputs, such as step functions [37]. This steady-state behavior is captured by the so-called DC gain and associated phase shift. In this limit, the frequency variable approaches zero, $\omega\to 0$. We have provided a general formulation for the DC gain in Methods IV.1.
@@ II.2 $\mathcal{H}_{2}$-Norm
In this work, the $\mathcal{H}_{2}$-norm serves as our principal analytical tool to quantify how the structure of a networked system shapes the amplification of inputs into outputs. For a given open system characterized by the triplet $A,B,C$, the $\mathcal{H}_{2}$-norm provides a universal and versatile measure of input-to-output amplification with general relevance in at least three fundamental scenarios: arbitrary time-varying inputs, sinusoidal inputs of arbitrary frequency, and noisy (stochastic) inputs. This classical optimal control performance index measures the total energy transfer from inputs to outputs [38, 39], quantifies variance amplification in stochastically driven linear systems [40], and coincides with the trace of the output controllability Gramian in linear time-invariant settings [38, 39]. Intuitively, it reflects both the energy of the impulse response [41] and the sensitivity of the system’s output to random input fluctuations.
The $\mathcal{H}_{2}$-norm squared for the above system measures the energy of the response to a unit impulse and can be computed as

### v2

@@ Abstract
Many biological, technological, and social systems can be effectively described as networks of interacting subsystems. Typically, these networks are not isolated objects, but interact with their environment through both signals and information that is received by specific nodes with an input function or released to the environment by other nodes with an output function. An important question is whether the structure of different networks, together with the particular selection of input and output nodes, is such that it favors the passing or blocking of such signals. For a given network and a given choice of the input and output nodes, the $\mathcal{H}_{2}$-norm provides a natural and general quantification of the extent to which input signals—whether deterministic or stochastic, periodic or arbitrary—are amplified. We analyze a diverse set of empirical networks and find that many naturally occurring systems—such as food webs, signaling pathways, and gene regulatory circuits—are structurally organized to enhance the passing of signals; in contrast, the structure of engineered systems like power grids appears to be intentionally designed to suppress signal propagation.
@@ I Introduction
Networked dynamical systems, composed of are omnipresent is nature and engineering, ranging from neuronal dynamics to electric power grids. They consist of many interacting units, each with its own intrinsic parameters and degrees of freedom, whose mutual coupling typically gives rise to rich collective behaviors. Usually, these networks are not isolated from their environment, but are subjected to external perturbations, operational changes, and external stimuli. Examples are neuron networks subject to sensory stimuli [1, 2], power grids that both absorb power produced at the generators and deliver power at the loads [3, 4, 5], and networks subject to the effects of environmental noise, such as ecological networks [6]. At the same time, networks produce outputs that affect the environment. Hence, the response of networked systems to external inputs essentially depends on the units chosen as input nodes and output nodes and the structure and characteristics of the network connectivity.
@@
While a large literature has studied control strategies for networks, see e.g., [7, 8, 9, 10], this paper focuses on the response of networked systems to various environmental inputs, not limited to control signals. References [11, 12, 13, 14, 15, 16] have studied the response of a networked system to external stimuli. These studies have primarily focused on graphs with symmetric connectivity or, at most, on asymmetric but normal adjacency matrices. Yet, most real-world networks are highly non-normal, strongly directed, and often exhibit pronounced hierarchical organization—structural features that profoundly shape their dynamical response [17, 18, 19, 20, 21, 22, 23, 24, 25]. In such networks, non-normality acts as an algebraic expression of directedness and emerges from strong asymmetries in the adjacency pattern with a clear source and sink structure, and induces strongly biased directions of flow. Structurally, highly non-normal networks are close to acyclic, with a dominant directed acyclic backbone and only sparse or weak cycles; this can be quantified, for example, by measuring the distance from the largest directed acyclic subgraph, a quantity that indicates how close the network is to a loopless structure [17, 18, 19].
Consistent with this view, recent work shows that increasing directedness and non-normality leads to flows that are more localized, strongly biased downstream, and effectively irreversible, so these architectures naturally favor one way signal transfer toward terminal components and sink nodes [22, 25]. Motivated by these properties, our approach does not assume symmetry and instead addresses the response to general input signals, such as sinusoidal, periodic, and stochastic, within a unified framework. Recent works have focused on how different nonlinearities affect the propagation patterns and stability in complex networks [26, 27]. Here, instead, we aim at characterizing whether specific types of networks tend to facilitate or block signal transmission, in terms of the combined effect of network structure as well as the corresponding dynamics.
@@
This paper examines how complex networks of coupled dynamical systems respond to environmental interactions, where external signals are received at designated input nodes and measured at specific output nodes [28, 29, 30, 31]. An open network is defined by (i) its internal topology i.e., the set $\mathcal{N}$ of the network nodes and the set $\mathcal{E}$ of directed, possibly weighted, edges that connect them, (ii) the set of input nodes $\mathcal{I}\subseteq\mathcal{N}$ where external stimuli are applied, and (iii) the set of output nodes $\mathcal{O}\subseteq\mathcal{N}$ where the response is measured. As illustrated schematically in Fig. 1, the network structure and its internal connectivity shape how input signals—including sinusoidal, periodic, and noisy—are transformed as they propagate from input to output, with different frequency components affected by amplification, attenuation, or distortion, depending on the structural properties of the network. Such signals are ubiquitous in natural and technological systems—for example, gene regulatory networks responding to circadian rhythms [32], power grids operating at standardized frequencies of 50–60 Hz [33], and neuronal circuits processing signals across a broad frequency range of roughly 0.1–100 Hz [34]—and we emphasize that our framework also applies to noise and arbitrary signals.
In panel (a) of Fig. 1, a low-frequency sinusoidal input may undergo phase shift and amplification, whereas a high-frequency input is attenuated. Panel (b) illustrates a network that suppresses noise. In panel (c), the output signal generally differs in shape from its corresponding input. In panel (d), the input contains both low- and high-frequency components; in the output, the low-frequency component passes through, while the high-frequency component is strongly attenuated and effectively filtered out. The Methods Section IV.2 presents a numerical example of signal transformation when it passes through a real connectome network.
@@
In this paper, we adopt a frequency-response viewpoint: we represent each system through its transfer function and quantify signal amplification using the $\mathcal{H}_{2}$-norm, which captures responses to sinusoidal, periodic, noisy, and more general inputs. Our analysis investigates how the topology of a network and the placement of its input and output nodes shape the amplification or attenuation of external signals. We establish a direct relationship between the network’s topology, its transfer function, and the controllability Gramian, which serves as a central tool in our study. This framework allows us to assess whether empirical networks tend to amplify (pass) or suppress (block) external inputs for given input-output configurations.
@@
Building on this framework, we next apply it to a broad collection of empirical networks from biology, technology, and social systems in order to uncover structural principles governing how they transform signals from input to output. Key findings are that many biological networks—such as food webs, signaling pathways, and gene regulatory networks—tend to exhibit architectures that enhance the passing of signals, supporting the efficient flow of biomass, information, or regulatory activity. This signal-passing behavior is especially evident in directed acyclic graphs (DAGs), namely directed networks that contain no directed cycles, which are known to closely approximate many natural real-world networks [18] and suggests that such architectures inherently promote signal propagation. Motivated by this, we analytically compute the DC gain i.e., the input to output gain when the input is a constant signal for DAG networks, thereby linking amplification or attenuation directly to the number and length of the input–output pathways. In contrast, engineered systems such as power grids—often modeled as mostly undirected (symmetric) graphs—appear to be purposefully designed to suppress signal amplification, reflecting the need for regulation of transmitted signals, such as voltage phase differences, to ensure stable and synchronized operation.
@@
FIG. 1: Networks as open systems. The schematic summarizes the goal of the paper in studying how an external signal is altered when it passes through a network. The network is subjected to various input signals from the environment, which are only received at the input nodes (in black.) The response of the network is measured at some output nodes (in orange). Several input signals that are considered in this paper and corresponding output signals are shown. In (a) a low-frequency sinusoidal input signal may experience a phase shift and amplification, while a high-frequency one may be damped. The case in (b) is that of a network that suppresses noise. In (c) the shape of a general output signal may not be the same as its corresponding input signal. In (d) the input signal contains both a low frequency and a high frequency components. In the output signal, the low frequency passes while the high frequency is strongly damped and thus effectively blocked.
@@ II.1 Frequency Analysis of Networks as Open Systems
Many natural and engineered systems—such as gene networks, power grids, and neural circuits—can be modeled as open networks, where each node evolves through intrinsic dynamics, interacts with neighbors, and communicates with the environment via inputs and outputs. While inherently nonlinear, such systems can often be approximated near equilibrium by a linear time-invariant (LTI) model, enabling powerful analysis through linear control theory.
Throughout this work we consider a general nonlinear model for the network dynamics of the form, $\dot{z}_{i}(t)=f(z_{i}(t))+\sigma\sum_{j}\tilde{A}_{ij}\,g(z_{i}(t),z_{j}(t))$, $i=1,..,N$, where $z_{i}(t)\in\mathbb{R}$ is the state of system $i$, the function $f:\mathbb{R}\rightarrow\mathbb{R}$ describes the intrinsic dynamics of each node, $g:\mathbb{R}\times\mathbb{R}\rightarrow\mathbb{R}$ the pairwise interactions, the scalar $\sigma$ is the global coupling strength, and $\tilde{A}\in\mathbb{R}^{N\times N}$ is the directed adjacency matrix of the network. With suitable choices of the functions $f$ and $g$, this general framework has been applied to model diverse networked systems, such as power grids, food webs with competitive and mutualistic interactions, gene-regulatory networks, and neuronal population dynamics. Since many real-world networked systems operate near a stable equilibrium, we study here the linearized dynamics around such an equilibrium, $\dot{x}_{i}=-A_{ii}x_{i}+\sum_{j}A_{ij}x_{j},i=1,\ldots,N,$ where the interdependencies among the linearized states $x_{i}$ of the network nodes are captured by the network Jacobian matrix $A=[A_{ij}]$. Since here we consider a stable fixed point, all eigenvalues of the Jacobian matrix $A$ have negative real part. For more details on the steps involved in the linearization, see the Methods Section IV.1. We remark that the network Jacobian $A$ shares a similar structure as the network adjacency matrix $\tilde{A}$, which is demonstrated with an example in the Methods Section IV.1.
This formulation leads to a networked system governed by the standard LTI equations:
@@
where $\boldsymbol{x}(t)\in\mathbb{R}^{N}$ is the network state vector, and $\boldsymbol{u}(t)\in\mathbb{R}^{m}$ is the vector of external input signals acting on the network, reflecting its open nature. The output vector $\boldsymbol{y}(t)\in\mathbb{R}^{p}$ collects signals from selected nodes. The matrix $A$ is the network Jacobian introduced before, the matrices $B$ and $C$ encode information on the nodes where inputs enter the network and the nodes from which outputs are measured, respectively. Each column of $B$ corresponds to one input signal and contains a single 1. If $B_{ij}=1$, this means that input $j$ is received at node $i$. Similarly, each row of $C$ corresponds to one output signal and also contains a single 1. If $C_{ij}=1$, this means that output $i$ is measured at node $j$. In short, each input is injected at exactly one node, and each output is read from exactly one node.
The LTI approximation provides a principled and analytically tractable framework for studying networked systems operating near stable fixed points, during short transients, or under weak perturbations, where linearized dynamics capture the dominant system behavior. This includes many biological and engineered systems functioning under steady-state conditions. Also systems near criticality, such as second-order phase transitions, where linear responses captures essential behavior [35], can be modeled as LTI systems. Even in synchronization problems, governing models in neurodynamics, the LTI framework remains valid when the system is close to the onset of collective oscillations, which typically arise through a Hopf bifurcation occurring at a specific mode (e.g., zero wavenumber or the zero coupling eigenvalue) [36, 37, 38]. In such cases, the dynamics can be linearized around the fixed point prior to bifurcation, yielding time-invariant Jacobians that characterize local behavior. Crucially, we find an important relation with the non-normality of the network Jacobian matrix, which is known to induce transient amplification and shape control energy[39] and is an ubiquitous feature of real-world networks. Finally, one should note that many-body interactions can also be accounted for within our framework. Indeed, when linearizing the dynamics around a stable fixed point, the interaction can, at leading order, be described by an effective pair-wise directed network through the Jacobian. In what follows, we restrict our attention to this first-order (linear) approximation, although higher-order terms e.g. in phase-reduction expansions can generate explicit many-body effective interactions beyond this description [40].
In this work, we adopt a frequency-domain approach to characterize the dynamics of networked open systems. Any periodic input can be decomposed into a superposition of sinusoids via Fourier analysis, enabling the study of the system’s frequency response to generic inputs. For the LTI model in Eq. (1), the transfer function [41] is defined as
@@
By evaluating the transfer function at $s=\jmath\omega$, where $\omega$ is the input frequency and $\jmath=\sqrt{-1}$, we probe the system’s steady-state response to sinusoidal inputs. After the transient has elapsed, the output oscillates at the same frequency $\omega$ of the input signal, but with altered magnitude and phase, given by
$\text{Magnitude}=20\log_{10}|G(\jmath\omega)|,\qquad\text{Phase}=\angle G(\jmath\omega).$ (3)
Specifically, if the input signal $u(t)=\sin(\omega t)$, then the output signal $y(t)=\text{Magnitude}\cdot\sin(\omega t+\text{Phase})$ [41]. The factor $20\log_{10}(\cdot)$ in Eq. (3) expresses the magnitude $|G(j\omega)|$ in decibels (dB), the conventional logarithmic unit for system gain. This frequency-based formulation underpins our analysis and enables principled insights into how the network structure shapes signal transmission and network performance.
An important case of the frequency response analysis arises in the low-frequency limit, which corresponds to the case of slowly varying or constant inputs, such as step functions [42]. This steady-state behavior is captured by the so-called DC gain. In this limit, the frequency variable approaches zero, $\omega\to 0$. We have provided a general formulation for the DC gain in Methods IV.3.
@@ II.2 $\mathcal{H}_{2}$-Norm
In this work, the $\mathcal{H}_{2}$-norm [43] serves as our principal analytical tool to quantify how the structure of a networked system shapes the amplification of inputs into outputs. For a given open network characterized by the triplet $A,B,C$, the $\mathcal{H}_{2}$-norm provides a universal and versatile measure of input-to-output amplification with general relevance in at least three fundamental scenarios: arbitrary time-varying inputs, sinusoidal inputs across the frequency spectrum, and noisy (stochastic) inputs. This classical optimal control performance index measures the total energy transfer from inputs to outputs [44, 43], quantifies variance amplification in stochastically driven linear systems [45], and coincides with the trace of the Gramian in linear time-invariant settings [44, 43]. Intuitively, it reflects both the energy of the impulse response [46] and the sensitivity of the system’s output to random input fluctuations.
The $\mathcal{H}_{2}$-norm squared for the system with transfer function (2) measures the energy of the response to a unit impulse and can be computed as
$\begin{split}\|G\|_{2}^{2}&:=\dfrac{1}{2\pi}\text{Tr}\left[\int_{-\infty}^{+\infty}G(\jmath\omega)^{H}G(\jmath\omega)d\omega\right]\\ &=\text{Tr}\left[\int_{0}^{+\infty}Ce^{At}BB^{\top}e^{A^{\top}t}C^{\top}dt\right]=\text{Tr}[CW_{c}C^{\top}],\end{split}$ (4)
where $W_{c}$ is the continuous-time controllability Gramian [47] and is defined as
$W_{c}=\int_{0}^{+\infty}e^{At}BB^{\top}e^{A^{\top}t}dt,$ (5)
which, for simplicity, we refer to as the Gramian throughout the paper. An important observation from Eq. (4) is that the $\mathcal{H}_{2}$-norm squared provides a measure of amplification to all frequencies between 0 and $\infty$, as well as to impulse time signals.
The matrix $W_{c}$ in (5) is symmetric and positive semi-definite and is finite if and only if the Jacobian matrix $A$ has all eigenvalues with negative real parts. The matrix $W_{c}$ can be easily calculated by solving the following continuous-time Lyapunov equation [47] for $W_{c}$:
$AW_{c}+W_{c}A^{\top}+BB^{\top}=0.$ (6)
For more information on the properties of the Controllability Gramian (5), see Methods IV.4, which in particular discusses the important modularity property of the trace of the output Gramian $\text{Tr}(W_{c}^{out})$. Namely, for any choice of the sets of input nodes and output nodes, the trace $\text{Tr}(W_{c}^{out})$ can be written as a sum of individual contributions, each one of which corresponds to a pair of input and output nodes (see Methods IV.4.) This has direct and important consequences, as it implies that if we assign the set of output nodes, a set of $k$ input nodes that minimizes (maximizes) $\text{Tr}(W_{c}^{out})$ is simply composed of the $k$ nodes that provide the largest (smallest) individual contributions. An analogous conclusion is obtained in the case that a given set of input nodes is assigned and one wants to select an optimal set of $k$ output nodes. Furthermore, Supplementary Note 1 relates the output standard deviation to the $\mathcal{H}_{2}$-norm via the Gramian spectrum, and Supplementary Note 2 provides upper bounds for the Bode integral in terms of the eigenvalues and eigenvectors of the matrix $A$.
Up to this point, we have examined the overall network amplification of input signals across frequencies, as quantified by the $\mathcal{H}_{2}$-norm. However, it is possible that a particular network will pass certain frequencies and block others, which can be studied by using Bode plots [48]. To illustrate the network response from a local, analytically tractable perspective, we next introduce a simple solvable model that clarifies how individual frequencies are processed and will serve as a reference point for our later analysis.
Example 1.
Any periodic input signal can be decomposed into different frequencies through its Fourier series decomposition, which allows the application of frequency response analysis to generic periodic input signals. We focus on the unidirectional chain network with $N$ nodes shown in Fig. 2, where we assume that node $1$ is the input node and node $N$ is the output node. For this example, the network Jacobian $A$ can be written as follows,